Main INDEX
Monthly INDEX
PREV
NEXT
User name lkaufman
Log entry time 10:49:16 on June10,2004
Entry number 125235
keyword=BPM/BCM calibration run info
A BPM/BCM calibration was taken during owl shift: HAPPEX DAQ run 1993.
There was no target in during this run.
The event numbers that I quote are only approximate of course:
Beam Condition Event numbers
--------------------------------------
no beam 1600-3400
50 uA 7200-9000
no beam 9700-14700
50 uA 15200-16700
no beam 17700-19500
40 uA 20500-22300
no beam 22900-24700
30 uA 25600-27400
no beam 28000-29800
20 uA 31200-33000
no beam 34000-35800
20 uA 36400-38200
no beam 39600-41400
10 uA 43400-45200
no beam 45900-47700
-------------------------------------
The BPM gains were set to forced gain.
bpm10 (iocse8): Gain X = 1175, Gain Y = 1140
bpm1,4a,4b (iocse10): Gain X = 1120, Gain Y = 1120
bpm8 (iocse9): Gain X = 1210, Gain Y = 1220
bpm12 (iocse9): Gain X = 890, Gain Y = 910
iocse9 is run in feedback mode, so the gains SHOULD be locked at these values, but there is no guarantee that they didn't relock.